« previous week | index | next week »
- 1:00pm PDT (20:00 UTC)
- Mozilla HQ, 1st floor conference table
- +1 866 432 7917 (US toll free)
- +1 334 309 0297 (International)
- join irc.mozilla.org #staffmeeting for attendance taking
- Ship shortly after FF2
- Take security fixes for FF2
- RC1 will be released this week
- We are already preparing for an RC2 - scheduled code freeze for this Fri
- driving queries available on Mozilla 1.8 Tinderbox page
- Focused on adding Vista Integration Hooks before Beta 1
- Circulating a proposed release schedule for TB2 to build, QA and l10n. Will go to a wider audience after we refine it based on available resources.
- Test Development
- Alice started working with Seneca College student Liz Chak - she will be participating in the windows performance testing effort
- Updated anti-phishing test cases [marcia]
- improved L10n test scrips. Added more cases, added results diffs to highlight changes in results across the 50 locales.
- Test Execution
- Ran top site spiker test for 1.8 for Windows, Mac, Linux [bc]
- Tested major/minor update on fx 2 [davel]
- smoke tests, BFTs, focused testing, L10n testing for FF2 RC1. Also started update testing over the weekend (thanks tomcat!) [marcia, tracy, timr, tomcat]
- Tested two sets of distros [tracy]
- Tools and Infrastucture
- Meeting with shaver, jonas, jst, davel, alice, bc, timr to discuss testing in the platform area. The main focus needed was getting the current tests running and the results more visible
- Logging - Redeveloped logging to better separate output logs by test target for easier public reporting of test results. Switched to gzip compression to ease web publishing. Made minor change to report output format to remove now redundant prefix. [bc]
- Community Test Day Sept 22nd. This was lead by Robcee. We had good turnout, about 15-20 people in the channel helping!
- Working with Seneca Collage on testing projects, including Win Perf. [robcee, alice]
- Bug Day Sept 19th [tracy and others]
- Mirror update
- Major mirror back online (@ 133% capacity)
- Additional mirrors coming online
- Seeking out more mirror capacity
- Bandwidth for FF2
- Looking for partners to help with burst capacity
- DB migrations successful, continued this week
- Rodan hardware failure
- planet.mozilla.org migrated to cluster due to hardware failure
- litmus, etc being brought up on a vm
- continuing to replace/remove old hardware
- Replacement hardware for landfill, corp giving.
- In conjunction with Seth & corp. giving, we are ordering a new server (running vmware) for community projects. We will host on a firewall'ed network, no corp/mozilla access.
- Continued work to improve telecommunications - keep the feedback coming.
Product Management and Marketing Updates
- Web site updates progressing.
- Press tour planning continues.
- Dria, Beltzner, and Mary attended the "An Event Apart" one-day web design symposium in Seattle. People were quite excited to see Mozilla there.
- Otherwise, no significant change from last week.
Webdev, Add-ons, AMO
- some work with build to test major updates
- Backend services, tests
- Work on getting public-facing pages ready for l10n tranlators, design teams
- Minor db changes, migration script changes
- mozilla.com l10n
- more meetings, clouserw or pkim has more info
- morgamic in Mt. View Wednesday-Friday
See the report on Mozilla Foundation activities for the week ending September 22, 2006. Note that Frank Hecker will miss today's call due to a family commitment.